OpenStudy (brinazarski):

In a particular human population, the allele frequencies for the A-B-O blood type locus are 0.2 (allele IA), 0.25 (allele IB), and 0.55 (allele i). a. What is the estimated frequency of individuals with blood type O?

OpenStudy (brinazarski):

Blood type A, actually...

OpenStudy (abb0t):

Are you familiar with hardy-weinberg? I think you can use that: p\(^2\)+2pq+q\(^2\)

OpenStudy (brinazarski):

I tried that... still coming up with the wrong answer; not sure why.

OpenStudy (abb0t):

Show me your process, I will let you know where you're going in the right or wrong direction. I am sure that you're doing it correct, maybe just some arthmetic issues. remember: p + q = 1 and if you have p\(^2\) = \(\sqrt{p^2}\) = p

OpenStudy (brinazarski):

Would that mean it's .2?

OpenStudy (brinazarski):

Cuz I did that with blood type O and it was .55 but that wasn't the answer... it was 0.3

OpenStudy (brinazarski):

And from that I assume it's looking for the squared number, but 0.04 isn't working in this case.

OpenStudy (brinazarski):

Because of that I'm hesitant to put .2 because I only have one try left.

OpenStudy (brinazarski):

Went to a help session and figured it out. Thank you though!